OpenStudy (anonymous):

Assume the sun is mostly composed of hydrogen atoms. Assume the mass of the sun is 1.7e30 kg. and the mass of a hydrogen atom is 1.672e-27 kg. how many atoms are there in the sun.

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

i did but it still says its wrong

OpenStudy (anonymous):

i am assuming that that is scientific notation and actually means \[1.7\times 10^{30}\] and \[1.672\times 10^{-27}\]

OpenStudy (anonymous):

so divide and subtract the exponents

OpenStudy (anonymous):

1.672 X 10^-27

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

\[\frac{1.7\times 10^{30}}{1.672\times 10^{-27}}\]

OpenStudy (anonymous):

that doesnt work tho!!!

OpenStudy (anonymous):

so you have two jobs: divide the numbers, i get \[1.0167\times 10^{30+27}=1.067\times 10^{57}\]

OpenStudy (anonymous):

divide the numbers out front \[1.7\div 1.672=1.067\] and then subtract the exponents \[30-(-27)=30+27=57\]

OpenStudy (anonymous):

its saying thats wrong